# Federated Learning with Reduced Information Leakage and Computation

### v1

@@ H.1 Details of Datasets
H.2 Details of the Model
Classifier and Loss function. In our experiment, we use logistic regression for Synthetic, Multilayer perceptron and CNN for Femnist, Stacked LSTM for Sent140. Note that without privacy concern, any classifier and loss function can be plugged in to Upcycled-FL. However, if we adopt objective perturbation as privacy protection, the loss function should also satisfy assumptions in Theorem C.2. Thus log loss $\mathscr{L}(z)=log(1+\exp(-z))$ is used. It’s not hard to verify that this loss function satisfies assumptions in Theorem C.2 by taking first and second order derivatives($\left|\mathscr{L}^{\prime}\right|\leq u_{1}=1\text{ and }\mathscr{L}^{\prime\prime}\leq u_{2}=\frac{1}{4}$). For Multilayer perceptron, we use a two layer network with hidden dimension 14*14. For CNN, a 3*3 convolutional layer with 3 channels is adopted, following by two linear layers(followed by activation and dropout) with dimension 1000 and 100. For Stacked LTSM, we use 2 layers with 128 dimension Sent140. The word embedding dimension of Sent140 is 300.
Implementation and Environment. Our code is implemented in Pytorch 1.11 [49]. We employ SGD as local optimizer, with momentum 0.5, and set number of epochs E to 20 at each iteration $m$. Learning rate is tuned to 0.05. We run the model on a single GTX 1080 Ti.
H.3 CNN on Femnist
Figure 4: Loss and accuracy on Femnist using CNN. Trained with fraction=0.3, straggler=90%. $\mu=0.1,\lambda=0.04$. Training(testing) accuracy is Upcycled-FL 96.36%(96.21%), FedProx 95.63%(95.74%).
H.4 Convergence on all datasets
H.4.1 90% Straggler
Figure 5: Loss and accuracy on synthetic datasets. Trained with fraction=0.3, straggler=90%. (a) Synthetic(iid); (b) Synthetic(0,0); (c) Synthetic(0.5,0.5); (d) Synthetic(1,1). $\mu=0.1,\lambda=0.04$ for Synthetic(iid), Synthetic(0,0), $\mu=1,\lambda=0.42$ for Synthetic(0.5,0.5), Synthetic(1,1).
(a) $\mu=0.1$
(b) $\mu=0.3$
(c) $\mu=0.5$
Figure 6: Loss and accuracy on Femnist. Trained with fraction=0.3, straggler=90%. $\lambda=0.04,0.12,0.21$ is set based on the value of $\mu=0.1,0.3,0.5$, respectively.
(a) $\mu=0.1$
(b) $\mu=0.3$
Figure 7: Loss and accuracy on Sent140. Trained with fraction=0.3, straggler=90%. $\mu=0.1,\lambda=0.04$ for (a), $\mu=0.3,\lambda=0.12$ for (b).
H.4.2 30% Straggler
Figure 8: Loss and accuracy on synthetic datasets. Trained with fraction=0.3, straggler=30%. (a) Synthetic(iid); (b) Synthetic(0,0); (c) Synthetic(0.5,0.5); (d) Synthetic(1,1). $\mu=0.1,\lambda=0.04$ for Synthetic(iid), Synthetic(0,0), $\mu=1,\lambda=0.42$ for Synthetic(0.5,0.5), Synthetic(1,1).
H.5 Privacy on all datasets
H.5.1 Objective perturbation
  • Convergence
    Figure 9: Objective perturbation on synthetic datasets. Trained with fraction=1, straggler=0%. $\mu=0.1,\lambda=0.04$. (a) Synthetic(iid); (b) Synthetic(0,0); (c) Synthetic(0.5,0.5); (d) Synthetic(1,1).
  • Impact of $\alpha$
    Figure 10: Impact of $\alpha$ on Femnist. Trained with fraction=1, straggler=0%. $\mu=0.1,\ \lambda=0.04$.
H.5.2 Output perturbation
Figure 11: Output perturbation on synthetic datasets. Trained with fraction=1, straggler=0%. $\mu=0.1,\lambda=0.04$. (a) Synthetic(iid); (b) Synthetic(0,0); (c) Synthetic(0.5,0.5); (d) Synthetic(1,1).
Figure 12: Objective and output perturbation on Femnist and Sent140. (a) shows the result using objective perturbation on Femnist. The average accuracy of private Upcycled-FL is 71.04%(70.71%), of private FedProx is 19.72%(18.82%). (b) shows the result using output perturbation on Femnist. The average accuracy of private Upcycled-FL is 79.43%(79.63%), of private FedProx is 78.43%(78.36%). (c) shows the result using objective perturbation on Sent140. The average accuracy of private Upcycled-FL is 74.36%(72.38%), of private FedProx is 69.94%(68.02%). (d) shows the result using output perturbation on Sent140. The average accuracy of private Upcycled-FL is 74.42%(72.52%), of private FedProx is 62.10%(62.13%).All trained with fraction=1, straggler=0%, $\mu=0.1,\lambda=0.04$. For objective perturbation, $\delta=0.001$.

### v2

@@ F.1 Details of Datasets
F.2 Details of Algorithm FedProx
Here we present the detailed algorithm of FedProx:
Algorithm 2 FedProx [Li et al., 2020]
1:  Input: $\mu>0$, $\{\mathcal{D}_{i}\}_{i\in\mathcal{I}}$, $\overline{\omega}^{0}$
2:  for $t=1$ to $T$ do
5:     Each client sends its local model to server.
$\overline{\omega}^{t+1}=\sum_{i\in\mathcal{I}}p_{i}\omega_{i}^{t+1}.$
7:  end for
F.3 Convergence on all datasets
Here we present the convergence and accuracy of the testing dataset in the final iteration for all datasets under 90% straggler and 30% straggler scenarios.
F.3.1 90% Straggler
(a) Syn(iid)
(b) Syn(0,0)
(c) Syn(0.5,0.5)
(d) Syn(1,1)
(e) FEMNIST
(f) Sent140
Figure 5: Convergence of Upcycled-FL and regular FL methods with the approximate same training time under 90% Straggler.
F.3.2 30% Straggler
We also conduct an experiment with a 30% straggler scenario to examine the convergence of our method. Table 3 demonstrates that our conclusions remain valid even when the straggler rate is relatively low.
Table 3: Average accuracy with 30% straggler on the testing dataset over four runs, the experimental setting is same as Table 1.
Method Dataset
Syn(iid) Syn(0,0) Syn(0.5,0.5) Syn(1,1) FEMNIST Sent140
FedAvg 98.50±0.18 80.30±1.01 82.60±0.72 78.81±2.10 83.06±1.30 74.29±0.12
Upcycled-FedAvg 99.30±0.25 80.77±1.09 83.09±0.14 80.74±2.71 84.41±0.09 75.18±0.50
FedAvgM 98.50±0.07 81.08±2.33 82.89±0.12 80.64±2.64 81.60±4.57 74.61±2.28
Upcycled-FedAvgM 99.12±0.17 82.20±1.80 82.68±0.19 80.43±2.51 82.31±3.17 74.16±2.84
FedProx 97.22±0.24 80.88±0.67 82.31±1.05 80.13±2.76 81.57±1.46 74.23±0.31
Upcycled-FedProx 98.24±0.21 81.52±0.82 83.05±0.14 81.73±0.84 82.85±1.60 74.83±0.14
Scaffold 98.28±0.14 79.95±1.67 79.54±0.72 72.27±4.68 76.85±5.55 75.76±0.21
Upcycled-Scaffold 99.52±0.18 79.46±1.62 81.83±2.08 73.34±1.13 77.69±2.39 77.01±0.30
FedDyn 98.17±0.08 82.06±0.61 80.97±1.04 78.65±3.83 84.39±0.95 75.88±0.11
Upcycled-FedDyn 98.57±0.22 82.08±0.68 82.80±1.42 81.11±2.64 85.34±2.47 75.90±0.33
pFedme 96.45±0.14 91.05±0.60 89.64±1.04 92.88±0.67 76.15±3.08 72.87±0.26
Upcycled-pFedme 96.89±0.27 91.06±0.61 89.40±1.13 92.64±0.90 77.36±3.59 73.48±0.49
FedYogi 99.38±0.28 81.06±2.53 80.65±1.32 79.14±1.63 79.98±7.83 77.63±0.29
Upcycled-FedYogi 99.60±0.14 81.77±2.06 81.72±1.02 80.32±0.76 81.96±4.28 77.78±0.85
F.3.3 30% Straggler
(a) Syn(iid)
(b) Syn(0,0)
(c) Syn(0.5,0.5)
(d) Syn(1,1)
(e) FEMNIST
(f) Sent140
Figure 6: Convergence of Upcycled-FL and regular methods with the approximate same training time under 30% Straggler.
F.4 Additional Privacy Experiments
In addition to privacy experiments on synthetic datasets, we also report the output perturbation and objective perturbation on real-world datasets: FEMNIST and Sent140 in Figure 7 and 8. Due to limited computational resources, we conduct these experiments using a subset of the previously established baselines.
Figure 7: Comparison of private Upcycled-FL and private FL methods using output perturbation (left) and objective perturbation (right) on FEMNIST.
Figure 8: Comparison of private Upcycled-FL and private FL methods using output perturbation (left) and objective perturbation (right) on Sent140.
F.5 Experiments on full-scale lowercase letter of FEMNIST
We conduct a convergence experiment with full-scale lowercase letters (26 classes) from the FEMNIST dataset, as shown in Table 4. We also conduct privacy experiments comparing our method with FedProx, as shown in Figure 9.
Table 4: Average accuracy on the large-scale FEMNIST dataset
Method FEMNIST(lowercase letters)
30% straggler 90% straggler
FedAvg $90.90\pm 4.55$ $88.50\pm 5.07$
Upcycled-FedAvg $91.26\pm 4.06$ $88.69\pm 4.94$
FedAvgM $91.02\pm 4.47$ $89.13\pm 4.35$
Upcycled-FedAvgM $91.05\pm 4.36$ $89.28\pm 4.05$
FedProx $91.91\pm 0.38$ $89.02\pm 4.53$
Upcycled-FedProx $92.08\pm 0.72$ $89.44\pm 4.30$
Scaffold $89.69\pm 3.92$ $89.11\pm 3.74$
Upcycled-Scaffold $89.77\pm 4.01$ $89.72\pm 3.90$
FedDyn $92.88\pm 0.30$ $91.42\pm 0.24$
Upcycled-FedDyn $93.56\pm 0.30$ $92.64\pm 0.37$
pFedme $81.36\pm 4.63$ $80.72\pm 1.47$
Upcycled-pFedme $85.62\pm 1.43$ $85.14\pm 0.93$
FedYogi $87.64\pm 1.05$ $85.61\pm 0.56$
Upcycled-FedYogi $88.89\pm 1.84$ $86.48\pm 1.30$
Figure 9: Comparison of private Upcycled-FedProx and private FedProx using output perturbation (left) and objective perturbation (right).
F.6 Training Time
We report the average training time to compare the communication cost in Table 5.
Table 5: Training time for output perturbation experiments on Syn(iid).
Time/s FedAvg FedAvgM FedProx Scaffold FedDyn pFedme FedYogi
Baseline $147.50\pm 6.69$ $147.47\pm 7.49$ $139.91\pm 10.81$ $153.88\pm 12.70$ $181.49\pm 17.60$ $467.42\pm 12.87$ $153.53\pm 13.72$
Upcycled $86.44\pm 9.23$ $84.95\pm 3.39$ $80.59\pm 3.29$ $95.09\pm 13.89$ $99.82\pm 3.32$ $259.54\pm 18.27$ $92.82\pm 2.24$